OpenStudy (anonymous):

Solve the system using elimination. 3X-9Y=18, X+2Y=1

OpenStudy (anonymous):

3X-9Y=18....(1) X+2Y=1......(2) multiply (2) by 3: 3X+6Y=3.....(3) ______________________ 3X-9Y=18....(1) - 3X+6Y=3.....(3) ......................... 0-15Y=15 Y=-1 sub in (2): X+2(-1)=1 X-2=1 X=3

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

@suckerofmath welcome